A biochemist is investigating the conversion of atmospheric nitrogen to unstable form by bacteria that inhabit the root system of certain legumes and needs to know the pressure in Kilo Pascal exerted by 1.25g of nitrogen gas in a flask of volume 500mL at 37°C.



1
Expert's answer
2021-02-25T03:42:13-0500

pV=nRT

p = nRT/V

n = m/M

M (N2) = 28 g/mol

R = 8.314 L kPa K−1 mol−1

n (N2) = 1.25/28 = 0.04 mol

p (N2) = (0.04 x 8.314 x (37 + 273))/ 0.5 = 206 kPa